OpenStudy (anonymous):

Mandy's Candies is most famous for their homemade fudge and homemade toffee. They make an average of 146 pounds of fudge and toffee combined each month. They sell a pound of fudge for $8.29 and a pound of toffee for $6.26. Mandy's averages $1,045.91 each month in fudge and toffee sales. On average, how many pounds each of fudge and toffee does Mandy's Candies sell in one month? 65 fudge; 81 toffee 54 fudge; 92 toffee 71 fudge; 75 toffee 68 fudge; 78 toffee

OpenStudy (danjs):

Looks like you need to set up a couple equations , and solve the system here

OpenStudy (danjs):

Lets define the variables first

OpenStudy (danjs):

Let F= Fudge T = Toffee

OpenStudy (danjs):

Total Pounds = #Fudge + #toffee total money = (price toffee)(#Toffee) + (Price fudge)(#fudge)

OpenStudy (danjs):

see the logic in those 2 statements

OpenStudy (danjs):

Using the variables that we defined T and F, putting those 2 statements into equations gives 146 = T + F (Pounds) 1045.91 = 6.26*T + 8.29*F (Money)

OpenStudy (danjs):

u understand that , or need more explanation

OpenStudy (anonymous):

A little more please bcuz I suck at math im only good at basics

OpenStudy (danjs):

do you see how i arrived at those 2 equations?

OpenStudy (anonymous):

yes

OpenStudy (danjs):

They make an average of 146 pounds of fudge and toffee combined each month Total = pounds toffee + pounds fudge 146 = T + F

OpenStudy (danjs):

They sell a pound of fudge for $8.29 and a pound of toffee for $6.26. Mandy's averages $1,045.91 each month in fudge and toffee sales. Total Sales = Price fudge x #pounds fudge sold + Price Toffee x # pounds toffee sold 1045.91 = 8.29*F + 6.26*T

OpenStudy (danjs):

The two equations are: 146 = T + F 1045.91 = 8.29*F + 6.26*T

OpenStudy (danjs):

good so far?

OpenStudy (anonymous):

yes

OpenStudy (danjs):

ok, to solve these and figure out which T and F values will satisfy both equations, we can do a substitution...

OpenStudy (danjs):

By that i mean, Solve the first equation for either T= or F=, and then substitute that value into the second equation, so you have only one variable in the equation.

OpenStudy (danjs):

146 = T + F 1045.91 = 8.29*F + 6.26*T Solving first for T: T = 146 - F Substitute that T value into the second equation: 1045.91 = 8.29*F + 6.26*(146 - F)

OpenStudy (danjs):

good on that part?

OpenStudy (anonymous):

yes

OpenStudy (danjs):

cool, now you have an equation with only one variable, F, we can solve that for F= 1045.91 = 8.29*F + 6.26*(146 - F)

OpenStudy (danjs):

First thing , distribute the 6.26 into the parenthesis by multiplying each by 6.26... 1045.91 = 8.29*F + 6.26*(146 - F) 1045.91 = 8.29*F + 913.96 - 6.26*F

OpenStudy (danjs):

Now, we can combine the like terms 1045.91 = 8.29*F + 913.96 - 6.26*F 1045.91 = (8.29-6.26)*F + 913.96 1045.91 = 2.03*F + 913.96

OpenStudy (danjs):

Subtract 913.96 from both sides 1045.91 - 913.96 = 2.03*F + 913.96 - 913.96 131.95 = 2.03*F

OpenStudy (danjs):

Divide both sides by 2.03 131.95 / 2.03 = (2.03/2.03) F 65 = F

OpenStudy (danjs):

There you have it, F = 65 , 65 pounds of Fudge was sold

OpenStudy (danjs):

good on that part?

OpenStudy (anonymous):

yes thank you for explaining it

OpenStudy (danjs):

welcome, That is all you need to do on this one, the only answer with F=65 is the first one. If you needed the pounds of toffee too, it is just one more simple step You have F=65 , put that into the first equation T+ F = 146 T + 65 = 146 T = 146 - 65 T = 81 pounds toffee